OpenStudy (anonymous):

Find the length of the diameter of a sphere with a surface area of 24.63 miles2.

OpenStudy (goldphenoix):

To find the surface area of a sphere, you use the formula: \[\large \large 4\pi r^2 = Surface Area\] Since we know that surface area of it, you can plug it in. \[\large \large 4\pi r^2 = 24.63 m^2\] We want to find the diameter of the sphere. To find it, you must find the radius. To find the radius, you must get rid of 4, pi, and the square. Let's get rid of the 4 first. To get rid of the 4, you must divide both by 4. 24.63 divided by 4 will give you about 6.16. Next you want to get rid of pi. To get rid of pi, you must divide both side by pi. 6.16 divided by pi will give you 1.96. Your equation should finally look like: \[\large \large r^2 = 1.96.\] We can get rid of the square by square rooting both side. The square root of 1.96 is 1.4. So your radius is: \[\large \large r=1.4\] You want to find the diameter. So you must time 1.4 by 2. Since radius time 2 will give you the diameter. 2 times 1.4 will give you 2.8. So the diameter is 2.8! I hope this help!

OpenStudy (anonymous):

Thanks a bunch